# FreeBlend: Advancing Concept Blending with Staged Feedback-Driven Interpolation Diffusion

### v2

@@ A.4 Details of Dataset Construction
Transports airplane, bicycle, car, ship
@@
The designed category list can be divided into four distinct groups based on their characteristics, as shown in Table 6. This categorization provides a clear organization of items based on their respective categories and reflects common concepts encountered in daily life.
A.5 Style Transfer Task
Figure 11 illustrates the impressive visual performance of our method on the style transfer task. Given a reference image and a content image, our training-free approach effortlessly transfers the content into the style of the reference, demonstrating both the adaptability and compatibility of our method for this task.
A.6 Details of Metrics
@@
where $\text{Score}(\cdot,\cdot)$ is the cosine similarity score computed by the CLIP model using the input image and prompt, $\text{Image}_{\text{blending}}$ represents the generated blended images, $\text{Image}_{\text{original}_{i}}$ denotes the original images of class $i$, and $n$ is the number of concepts blended. $\text{Prompt}_{i}$ refers to the prompt for class $i$, formatted as “a photo of a {class}” to minimize noise in the image generation process. This metric can also be interpreted as the high-dimensional distance between the blended concepts and the original ones. It represents the total distance between the blended concept and each of the original concepts. A larger score generally indicates the dimensional center of the original concepts.
DINO-BS (Liu et al., 2025). Grounding DINO is an open-set object detection model designed to improve the ability to interpret human language inputs through tight modality fusion and large-scale grounded pre-training. The model excels at detecting unseen objects and can perform effective zero-shot transfer learning, even for categories not encountered during training, by leveraging language prompts. To guide the model in detecting blended objects, we designed two specific prompts: “a {concept_a} with blending features from {concept_b}” and “a {concept_b} with blending features from {concept_a}”. These prompts help the model appropriately recognize and identify the blended objects.
CLIP-IQA (Wang et al., 2023). This metric employs a thoughtfully crafted prompt pairing strategy to reduce linguistic ambiguity and effectively harness CLIP’s pretrained knowledge for assessing visual perception. In our task, we design the specific prompt pairs like (“mixed”, “dull”), (“blending features from two different objects”, “natural object from one object”).
HPS (Wu et al., 2023). This metric more accurately predicts human preferences for generated images. The designed prompt is “a photo of a blended object combining mixed features from {concept_a} and {concept_b}”.
A.7 Baseline Methods
@@
  • MagicMix (Liew et al., 2022). The original MagicMix method modifies the input image using a text prompt from a different class to transfer the image’s characteristics. In our implementation, we first generate the original image using a class text prompt with Stable Diffusion then pass it into the original MagicMix pipeline.
  • Composable Diffusion (Liu et al., 2022). This method decomposes the text description into components, each processed by a different encoder to produce a latent vector capturing its semantic information. Multiple diffusion models are then used to independently generate images for each component. The Conjunction operation combines these components into a single image. This approach excels in zero-shot compositional generation, allowing the model to create novel combinations but not blending of objects and their relationships, even without prior training on such combinations.
  • unCLIP-T (Wang et al., 2023). The unCLIP model leverages images as conditioning inputs, replacing traditional text prompts, by training several linear layer adapters. In this implementation, image embeddings are treated similarly to text embeddings, with their average embedding being computed to serve as the conditions.
  • unCLIP-U (Wang et al., 2023). This implementation is akin to unCLIP-T, with the key difference being that the image embeddings are integrated using the UNET method.
  • TEXTUAL (Melzi et al., 2023). This method encodes the text prompts into text embeddings and computes their average embedding to input into Stable Diffusion for generating blended images. However, it is unstable within the embedding space and constrained by the semantic scope of the concepts involved.
  • UNET (Olearo et al., 2024). This method uses the first text embedding for the conditions of downsampling of U-Net (Ronneberger et al., 2015) and the second one for the upsampling of the U-Net.
A.8 User Study
The screenshots in Figure 12 display our questionnaires used to assess the validation of different methods.
A.9 More Analysis of Blending Results
@@
A.9.1 Types of Blending Concept Pairs
Biological Blending. This type of blending involves the blending of entities from the natural world. For example, when animals with similar characteristics, such as cats and dogs, are blended, the resulting output tends to exhibit a high degree of coherence and natural harmony. In contrast, blending animals with plants often leads to situations where the plant components are relegated to background elements, rather than being seamlessly integrated into the final composition. This phenomenon suggests that intra-category blending (e.g., animal-animal or plant-plant) typically yields more naturalistic and fluid results, whereas inter-category blending may result in less cohesive outputs.
Decorative Object-Feature Blending. This form of blending involves the decoration of objects with specific attributes or features. For instance, blending a car with a rose may yield an aesthetic result in the form of a floral-patterned vehicle. Similarly, blending a tomato with a bicycle could result in a red-painted bicycle. In contrast, combining a lily with a bicycle may produce a bicycle adorned with floral elements or set against a backdrop evoking a garden. This type of blending is particularly notable for its capacity to generate creative reinterpretations of objects, often yielding visually appealing and conceptually intriguing combinations.
Material and Pattern Blending. This type refers to the integration of materials—such as stripes, textures, or other patterns—into various objects. The primary characteristic of this form of blending lies in the inherent compatibility of materials with a wide range of object types, thereby facilitating smoother transitions and more seamless integrations across different categories. This suggests that materials have universal applicability, enabling them to blend effectively with diverse objects regardless of their inherent features.
Inherently Unrelated Blending. This form of blending presents significant challenges. When attempting to blend entities from vastly different categories, such as a phone with a dinosaur or an airplane with a dog, the resulting combinations often appear disjointed or nonsensical. This highlights the difficulty in merging categories that exhibit substantial differences in their fundamental characteristics, suggesting that certain combinations may be inherently incompatible.
Broad and General Blending. This type can also complicate the blending process. For example, blending a concept as broad as ”human” tends to result in less coherent outputs, likely due to the wide range of attributes and features that fall under such a generalized category. In contrast, simple, well-defined descriptors such as ”black” tend to blend more effectively, as they are associated with clear, unambiguous characteristics that leave little room for interpretive variation.
Contextual and Background Blending. Context plays a critical role in the outcome of certain blending processes. Many concepts inherently carry with them specific contextual or environmental associations. For example, when the concept of a ”squirrel” is blended, it is often accompanied by a contextual backdrop of trees or forests. These environmental associations are embedded within the very concept of the entity itself, influencing the final blending output and shaping how the resulting image or combination is perceived.
A.9.2 Factors of Blending Concepts
Subject and Background. The effect of concept blending largely depends on the categories to which the blended concepts belong—whether they are subjects or backgrounds. When two concepts, such as a dog and a cat, both belong to the subject category, they are typically blended into a new subject, or one of them may dominate the resulting blend. In contrast, when the blend involves a subject and a background, such as a sheep and a cactus, the result may depict the sheep in a natural environment with cactuses, or the sheep may take on cactus-like characteristics, such as cactus-inspired wool. It is important to note that the identity of a concept can shift depending on how it is paired. For instance, while a cactus serves as the background concept in the pairing ”cactus, sheep,” it becomes part of the subject pair when blended with a rose, as in ”cactus, rose.” Thus, the role of a concept—whether as a subject or background—can significantly alter its representation in the blending process.
@@ A.9.2 Factors of Blending Concepts
Distance between Concepts. Distance is another crucial factor that determines the effectiveness of the blending. This principle is illustrated through t-SNE analysis in Section A.1, which demonstrates how the proximity of two concepts within an embedding space influences their ability to blend. When two concepts are close to each other in this space, they blend easily, forming a cohesive blend. However, as the distance between the concepts increases, the blending becomes less seamless. For example, concepts that are close in proximity, such as cats and dogs, tend to blend effortlessly, with the final output lying somewhere between both concepts. When concepts are moderately distant, such as a cactus and a rose, partial blending occurs, with the resulting image leaning more towards one concept while still maintaining some elements of the other. As the distance between concepts becomes even greater, such as in the case of an airplane and a bird, the blending becomes limited, and the final result often favors one concept, with the other remaining a secondary influence.
Special Cases in blending. Certain concepts possess intrinsic backgrounds or properties that affect blending. For instance, objects like tables and chairs, which are complementary in nature, are more likely to be presented together within the same space rather than being blended into a single form. Similarly, the blending of elements like water and fire may not result in a fluid combination but instead transform into something entirely different, like stone (result from TEXTUAL method), as the inherent properties of the elements conflict with each other.
A.10 Limitations and Future work
Our method has several limitations that negatively affect the blending results. First, the latent interpolation calculation is unstable during generation, often leading to chaotic or overlapping outcomes. Second, when blending multiple concepts (e.g., three or four), the results are unsatisfactory, as the UNET method is not designed to handle more than two conditions simultaneously. Third, the generation process is prone to fluctuation, lacking consistency and stability. Fourth, the Stable Diffusion model exhibits a bias towards different concepts, necessitating the adjustment of the parameter $\gamma$, which is cumbersome and inefficient. Therefore, we aim to explore more stable, cost-effective, and training-free generation methods, as well as seek a more structured approach for blending multiple concepts.
A.11 More Visual Results
Figure 13 and Figure 14 present several of our blending results, highlighting the strong capabilities of our method for both concept blending and generation tasks.
@@ A.11 More Visual Results
car-brocolli car-corgi car-lemon car-orange car-teddy
@@
eagle-cat eagle-lemon eagle-orange eagle-parrot eagle-penguin
@@
knight-lemon knight-lion knight-plane knight-panda knight-robot
clock-cactus clock-eagle clock-helicopter clock-ice cream clock-lemon
@@
burger-coca cola burger-dog burger-ice cream burger-robot burger-strawberry

### v3

@@ Appendix D Details of Dataset Construction
Transport airplane, bicycle, car, ship
@@
The designed category list can be divided into four distinct groups based on their characteristics, as shown in Table˜8. This categorization provides a clear organization of items based on their respective categories and reflects common concepts encountered in daily life.
Appendix E Why Vanilla Stable Diffusion Fails
Figure 9: Relying solely on prompts for concept blending, we find that Stable Diffusion 2.1 often fails to effectively merge the concepts as instructed. Instead, it typically depicts the two concepts as separate, coexisting elements within the image.
Appendix F More Style Tasks
Figure 10: When compared to other image interpolation methods, we find that AID [qiyuan2024aid] exhibits a stronger attention bias towards maintaining the quality of generated images, but it does not effectively blend concepts together. DiffMorpher [zhang2024diffmorpher] is designed to smoothly transfer one image to another, but when applied to unrelated concepts, it generates white noise. On the other hand, Image Interpo [wang2023interpolating] interpolates between two different images but suffers from significant overlapping issues.
Image Interpolation Task. Figure˜10 presents a qualitative comparison with other image interpolation methods, demonstrating that our method achieves the best results.
Figure 11: Comparison with the setting of MagicMix [liew2022magicmix]. Given a reference image and a prompt, the model can transfer the reference image to another concept. From the results, we observe that when there is a significant semantic and shape gap between the original and target concepts, MagicMix struggles to modify the original image while preserving its features, without introducing new ones. This highlights that our method is capable of handling concepts with substantial semantic and shape differences.
Switching One Condition from Image to Text. Figure˜11 presents the qualitative comparison between MagicMix [liew2022magicmix] and our method.
Style Transfer Task. Figure˜12 illustrates the impressive visual performance of our method on the style transfer task. Given a reference image and a content image, our training-free approach effortlessly transfers the content into the style of the reference, demonstrating both the adaptability and compatibility of our method for this task.
Figure 12: Adapting our method to the style transfer task.
Appendix G Details of Metrics
@@
where $\text{Score}(\cdot,\cdot)$ is the cosine similarity score computed by the CLIP model using the input image and prompt, $\text{Image}_{\text{blending}}$ represents the generated blended images, $\text{Image}_{\text{original}_{i}}$ denotes the original images of class $i$, and $n$ is the number of concepts blended. $\text{Prompt}_{i}$ refers to the prompt for class $i$, formatted as “a photo of a {class}” to minimize noise in the image generation process. This metric can also be interpreted as the high-dimensional distance between the blended concepts and the original ones. It represents the total distance between the blended concept and each of the original concepts. A larger score generally indicates the dimensional center of the original concepts. Note that irrelevant content could potentially receive a high score using this metric due to unrelated features. However, this scenario is unlikely to occur, as our generated images are based on relevant concepts.
DINO-BS [liu2025grounding]. Grounding DINO is an open-set object detection model designed to improve the ability to interpret human language inputs through tight modality fusion and large-scale grounded pre-training. The model excels at detecting unseen objects and can perform effective zero-shot transfer learning, even for categories not encountered during training, by leveraging language prompts. To guide the model in detecting blended objects, we designed two specific prompts: “a {concept_a} with blending features from {concept_b}” and “a {concept_b} with blending features from {concept_a}”. These prompts help the model appropriately recognize and identify the blended objects.
CLIP-IQA [wang2023exploring]. This metric employs a thoughtfully crafted prompt pairing strategy to reduce linguistic ambiguity and effectively harness CLIP’s pretrained knowledge for assessing visual perception. In our task, we design specific prompt pairs like (“mixed”, “dull”), (“blending features from two different objects”, “natural object from one object”).
HPS [wu2023human]. This metric more accurately predicts human preferences for generated images. The designed prompt is “a photo of a blended object combining mixed features from {concept_a} and {concept_b}”.
Appendix H Baseline Methods
@@
  • MagicMix [liew2022magicmix]. The original MagicMix method modifies the input image using a text prompt from a different class to transfer the image’s characteristics. In our implementation, we first generate the original image using a class text prompt with Stable Diffusion, and then pass it into the original MagicMix pipeline.
  • Composable Diffusion [liu2022compositional]. This method decomposes the text description into components, each processed by a different encoder to produce a latent vector capturing its semantic information. Multiple diffusion models are then used to independently generate images for each component. The Conjunction operation combines these components into a single image. This approach excels in zero-shot compositional generation, allowing the model to create novel combinations but not blending of objects and their relationships, even without prior training on such combinations.
  • unCLIP-T [wang2023exploring]. The unCLIP model leverages images as conditioning inputs, replacing traditional text prompts, by training several linear layer adapters. In this implementation, image embeddings are treated similarly to text embeddings, with their average embedding being computed to serve as the conditions.
  • unCLIP-U [wang2023exploring]. This implementation is akin to unCLIP-T, with the key difference being that the image embeddings are integrated using the UNET method.
  • TEXTUAL [melzi2023does]. This method encodes the text prompts into text embeddings and computes their average embedding to input into Stable Diffusion for generating blended images. However, it is unstable within the embedding space and constrained by the semantic scope of the concepts involved.
  • UNET [olearo2024blend]. This method uses the first text embedding for the conditions of downsampling of U-Net [ronneberger2015u] and the second one for the upsampling of the U-Net.
  • AID [qiyuan2024aid]. This paper introduces Attention Interpolation via Diffusion (AID), a training-free method for improving conditional image interpolation in text-to-image models. AID enhances consistency and smoothness, with a variant (PAID) enabling user-guided interpolation. Experiments show AID outperforms traditional methods and enhances control for image editing.
Appendix I User Study
Figure 13: Screenshots of user study.
The screenshots in Figure˜13 display our questionnaires used to assess the validation of different methods.
Appendix J More Visual Results
Figure˜14 and Figure˜15 present several of our blending results, highlighting the strong capabilities of our method for both concept blending and generation tasks.
Appendix K More Analyses of Blending Results
@@ Appendix I User Study
K.1 Results from Different Categories of Concepts
Table 9: Mean scores for different categories.
Category CLIP-BS($\uparrow$) DINO-BS($\uparrow$) CLIP-IQA($\uparrow$) HPS($\uparrow$)
Animals 9.4410 0.2460 0.5160 0.2800
Common objects 7.9810 0.4160 0.5270 0.3000
Nature 9.2250 0.4190 0.3330 0.2760
Transports 9.3100 0.2280 0.5430 0.3060
Table 10: Mean scores for blending category pairs.
Category Pair CLIP-BS($\uparrow$) DINO-BS($\uparrow$) CLIP-IQA($\uparrow$) HPS($\uparrow$)
Animals & Common 8.7110 0.3308 0.5212 0.2908
Animals & Nature 9.3330 0.3323 0.4244 0.2780
Animals & Transports 9.3760 0.2369 0.5294 0.2930
Common & Nature 8.6030 0.4170 0.4298 0.2880
Common & Transports 8.6450 0.3217 0.5348 0.3030
Nature & Transports 9.2680 0.3232 0.4380 0.2910
Table˜9 and Table˜10 show the blending scores of four metrics for different category pairs. The CLIP-BS in Table˜9 indicates that similar types of objects, such as animals, nature, and transport, achieve higher scores. However, for common objects, the score is lower due to the larger diversity within this category. Regarding DINO-BS, categories in nature tend to blend more successfully. In the case of HPS, we observe that transport categories generate higher-quality images. In Table˜10, the pattern is similar to that in Table˜9, where the best scores align with the original categories. Additionally, compared to nature categories, animals tend to blend better, which we hypothesize is due to biological reasons discussed below.
K.2 Factors of Blending Concepts
@@
Distance between Concepts. Distance is a crucial factor that determines the effectiveness of the blending. This principle is illustrated through t-SNE analysis in Appendix˜A, which demonstrates how the proximity of two concepts within an embedding space influences their ability to blend. When two concepts are close to each other in this space, they blend easily, forming a cohesive blend. However, as the distance between the concepts increases, the blending becomes less seamless. For example, concepts that are close in proximity, such as cats and dogs, tend to blend effortlessly, with the final output lying somewhere between both concepts. When concepts are moderately distant, such as a cactus and a rose, partial blending occurs, with the resulting image leaning more towards one concept while still maintaining some elements of the other. As the distance between concepts becomes even greater, such as in the case of an airplane and a bird, the blending becomes limited, and the final result often favors one concept, with the other remaining a secondary influence.
Appendix L Limitations and Future Work
In the case of the Stable Diffusion model, the bias towards different concepts requires adjusting the parameter $\gamma$, a process that is cumbersome and inefficient. Moreover, when blending three or more concepts, it becomes challenging because the UNET architecture can only support two conditions. Therefore, our goal is to explore more stable, cost-effective, and training-free generation methods, while also seeking a more structured approach for blending multiple concepts.
Appendix M Impact Statement
This paper presents a contribution aimed at advancing the field of concept blending. We explicitly state that our work is intended for academic research and commercial applications, provided such uses are authorized by the author. However, we strongly emphasize that we do not condone, nor do we support, the use of our research to generate harmful, unethical, or unlawful content. The generation of unsafe or offensive images that violate ethical standards, legal regulations, or societal norms is strictly prohibited. We use the Safe Checker from Stable Diffusion [rombach2022high] to prevent the generation of unsafe content. As researchers, we believe it is our responsibility to ensure that our work is used for the betterment of society and in a manner that aligns with fundamental ethical principles.
@@
car-brocolli car-dog chicken-donkey super car-orange car-teddy
@@
eagle-cat eagle-lemon eagle-orange eagle-parrot dog-plane
knight-lemon knight-lion knight-plane knight-panda knight-robot
@@
clock-cactus orange-car cat-phone clock-ice cream lily-lemon
@@
city-monument burger-dog burger-ice cream ship-lamp burger-strawberry
dog-Roman statue cathedral-fireworks pizza-tiramisu cake dog-astronaut cat-lavender
red ferrari-hello kitty woman-opera player teddy bear-girl car-clown anime-Renaissance